A is a solution of hydrochloric acid containing 3.60g of Hcl in 1dm³ of solution. B is a solution of impure of sodium carbonate.25cm³ of B require 22.7cm³ of A for complete reaction.Given that 5.00g of impure B was dissolved in 1dm³ solution,what is the percentage by mass of pure sodium carbonate in the impure sample​


1
Expert's answer
2021-10-25T03:46:48-0400

Solution:

The molar mass of HCl is 36.46 g/mol

The molar mass of Na2CO3 is 105.99 g/mol


Calculate the molarity of solution A:

(3.60 g HCl) × (1 mol HCl / 36.46 g HCl) × (1 / 1dm3) × (1 dm3 / 1 L) = 0.09874 mol HCl dm-3

CM(HCl) = 0.09874 mol/dm3


Balanced chemical equation:

Na2CO3(aq) + 2HCl(aq) → 2NaCl(aq) + H2O(l) + CO2(g)

According to stoichiometry:

n(Na2CO3) = n(HCl) / 2

2 × CM(Na2CO3) × V(Na2CO3) = CM(HCl) × V(HCl)

2 × CM(Na2CO3) × (25 cm3) = (0.09874 mol/dm3) × (2 × 22.7 cm3)

CM(Na2CO3) = (0.09874 mol/dm3 × 22.7 cm3) / (2 × 25 cm3) = 0.04483 mol/dm3

CM(Na2CO3) = 0.04483 mol/dm3


Calculate the moles of Na2CO3:

(0.04483 mol Na2CO3 / 1 dm3) × (1 dm3) = 0.04483 mol Na2CO3


Calculate the mass of Na2CO3:

(0.04483 mol Na2CO3) × (105.99 g Na2CO3 / 1 mol Na2CO3) = 4.75 g Na2CO3

Mass of pure Na2CO3 is 4.75 grams


Calculate the percentage by mass of pure Na2CO3 in the impure sample:

%Na2CO3 = (Mass of pure Na2CO3 / Mass of impure sample) × 100%

Mass of impure sample is 5.00 grams

Therefore,

%Na2CO3 = (4.75 g / 5.00 g) × 100% = 95%

%Na2CO3 = 95%


Answer: The percentage by mass of pure Na2CO3 in the impure sample​ is 95%